OpenStudy (anonymous):

Find the consumer's surplus if the demand for an item is given by D(x)=72-x2, assuming supply and demand are in equilibrium at x=6

OpenStudy (amistre64):

consumer surplus is everything above the equilibrium line; but whats the equation for supply?

OpenStudy (amistre64):

helps if I read the qwestun all the way to the end :)

OpenStudy (amistre64):

is that x^2?

OpenStudy (amistre64):

its the area under the curve..... i see that much :)

OpenStudy (amistre64):

y = 36 when x = 6; so we need the area under the curve from [0,6] of the equation 36-x^2

OpenStudy (anonymous):

so how do we find the area

OpenStudy (amistre64):

36x - (x^3)/3; at x=0 we get 0 at x = 6 we get... 36(6) - (6^3)/3 as the ansswer

OpenStudy (amistre64):

its taking the antiderivative/indegral of the equation with the bounds established...

OpenStudy (anonymous):

Ohhh ok thanksx:D

OpenStudy (amistre64):

I get 144 as my answer

OpenStudy (anonymous):

OK i got 2268??

OpenStudy (amistre64):

at x = 6; the y(equilib line) = 36 right?

OpenStudy (amistre64):

so we just want the area under the top part for our consumer surplus

OpenStudy (amistre64):

drop your equation by 36 to get it to the x axis: 72-36 -x^2 = 36 - x^2 integrate 36 - x^2 to get; 36x - (x^3)/3; and solve for x at 0 and 6

OpenStudy (anonymous):

ook i know wat i did wrong

OpenStudy (amistre64):

if you dont drop it; then you have to subtract the bottom part :)

OpenStudy (anonymous):

gotcha! :D

OpenStudy (amistre64):

:) this is the graphical representation ;)

OpenStudy (anonymous):

ok thanx